OpenStudy (anonymous):

how to graph Arg(Z)<=pi/3 where Z is complex

OpenStudy (anonymous):

remember while graphing any inequality u first draw the graph of the equation

OpenStudy (anonymous):

so first draw arg x=pi/3

OpenStudy (anonymous):

so let Z=x+yi .: tan(pi/3)=y/x .:y=\[\sqrt{3} x \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hm next?

OpenStudy (anonymous):

the region below this line corresponds to arg<pi/3

OpenStudy (anonymous):

its Arg(Z) -pi<Arg(Z)<=pi so?

OpenStudy (anonymous):

below this line till the x-axis

OpenStudy (anonymous):

how u know

OpenStudy (anonymous):

cz that region means the arg is less than pi/3